OpenStudy (anonymous):

(3x+6)^2= Rewrite without parentheses and simplify!! PLSS

OpenStudy (anonymous):

(3x+6)(3x+6) 3x*3x = 9x^2 3x*6 = 18x 6*3x = 18x 6*6 = 36 Add all these together. 9x^2 + 18x + 18x + 36 9x^2 = 36x + 36 Hope this helps.

OpenStudy (anonymous):

you are amazing!!!!

OpenStudy (anonymous):

Is there another one you want to work on?

OpenStudy (anonymous):

no thank you for now im good. i have another q if ud like :-)